AisanDoll:

solve for x y=2x y=20 Simplify your answer as much as possible.

carmelle:

Are you familiar with solving for x? Do you know what the first step is?

AisanDoll:

no

singlelilmomma:

20=2x 20/2=x X=10

AisanDoll:

singlelilmomma:

x=10 Step-by-step explanation: 20 divided by 2 equals 10.

carmelle:

Thats ok :) I'll help you we already know wha our `y-value` is so we can replace y with `20` so we have: `20 = 2x` To find the value of x, we will have to divide both sides by 2 \[\frac{ 20 }{ 2 } =~ ?~ \frac{ 2x }{ 2 } = ?\]

singlelilmomma:

i already helped her

AisanDoll:

thx

singlelilmomma:

no problem

carmelle:

do you understand how to answer the question though?

AisanDoll:

AisanDoll:

8

carmelle:

you need help with question 8?

AisanDoll:

5,8,9

singlelilmomma:

its 10 i gave you the answer already

AisanDoll:

@aisandoll wrote:
2

singlelilmomma:

Use the graph to determine which statement is true about the end behavior of f(x). help me

carmelle:

@singlelilmomma wrote:
Use the graph to determine which statement is true about the end behavior of f(x). help me
post your own question so that someone can help :)

carmelle:

I cant see the question on the paper, the picture is blurry @aisandoll

singlelilmomma:

its for edu

carmelle:

ok lets start with 5, is the equation 2 × 8 - 4 - 16 + 4

AisanDoll:

2x8 =16 16/4=4

AisanDoll:

2x8-4-16/4 16-4-16/4 16-4-4 12-4 8

carmelle:

Yes that is correct!! Great job!! do you know what the answer is for question 6?

carmelle:

or would you like me to help?

AisanDoll:

question 8

AisanDoll:

and 9

carmelle:

ok, have you learned absolute values before?

AisanDoll:

no

AisanDoll:

8 is |-5| and |14|

AisanDoll:

question 8

carmelle:

Here is an example: |-10| = 10 |10| = 10 another one if you dont understand... |-15| = 15 |15| = 15 based on what I said do you understand it now? The absolute value is the` distance between a number and zero.`

AisanDoll:

thx

carmelle:

wait.. do you know what the answers would be then?

AisanDoll:

ye

carmelle:

what did you get as your final answer?

AisanDoll:

5 and 14

carmelle:

Thats correct!!

carmelle:

For question 9, we need to get w by itself so we have to add 4.8 on both sides. `w - 4.8 = 9.2` -4.8 + 4.8 = ? 9.2 + 4.8 = ?